OpenStudy (anonymous):

A baseball is hit at 40 m/s at an angle of 30 degrees to the horizontal. a) How high will it go? b) When will it reach that height? c) What will be its horizontal distance from the batter at that time?

OpenStudy (anonymous):

Someone please help me!

OpenStudy (anonymous):

I'm not too into the physics course, our class just started

OpenStudy (anonymous):

time to reach the maximum height use Vf= Vy +gt 0=20 m/s -9.83t solve to find t Vy=v sin 30 (the vertical component of th initial velocity

OpenStudy (anonymous):

how high will it go? use Vf squared =Vy squared + 2 gh at max height Vf=0 ; Vy is vertical component of initial velocity v 0=40 sin 30 (squared) + 2(-9.83) x h 19.66h= 400 solve to find h

OpenStudy (anonymous):

Range= Vx x time of flight Vx = Horizontal component of initial velocity=v x Cos 30 time of flight = 2 x time to reach maximum height hope this has been of some help

OpenStudy (anonymous):

you are given the initial velocity of 40m/s and the angle \[\theta = 30\]. this is a projectile motion when you reach the top the final velocity will zero \[v =0\]. the maximum height that you are looking is the vertical component(y-axis) of the motion then the initial velocity in the y is then\[v _{0}\sin \theta = 40\sin 30 = 20m/s\]. then i would use the equation \[v^{2} = u ^{2} + 2as\], where \[a = g = -9.8 m/s ^{?}\] is the accelaration due to gravity acting downwards and \[s\] is the height you looking for.\[v^{2} = u ^{2} + 2as \rightarrow s = \frac{ 200 }{ 9.8} m\]